OpenStudy (anonymous):

find the exact value cos^-1 [tan(pi/4)]

OpenStudy (anonymous):

A. -1 B. 0 C. 1/2 D. 1

OpenStudy (anonymous):

do you know what \(\tan(\frac{\pi}{4})\) is ?

OpenStudy (anonymous):

To convert radians to degrees, multiply by \(\frac{ 180 }{ \pi }\) So: \[\frac{ \pi }{ 4 } \times \frac{ 180 }{ \pi }\]

OpenStudy (anonymous):

sooo im going with c?

OpenStudy (anonymous):

\[\frac{ \pi }{ 4 } \times \frac{ 180 }{ \pi } = 45\] SO: \[\tan( \frac{ \pi }{ 4 } ) = \tan(45)\]

OpenStudy (anonymous):

not really clear why you need to convert these functions are functions of numbers, not "degrees" \[\tan(\frac{\pi}{4})=1\]

OpenStudy (anonymous):

well yeah, but at the end its the same thing lol :P

OpenStudy (anonymous):

\[\tan(\frac{ \pi }{ 4 }) = \tan(45) = 1\]

OpenStudy (anonymous):

not really trig functions are functions of numbers as such, the numbers correspond to angles, but only if measured in radians

OpenStudy (anonymous):

\[\cos^-1 [\tan(\frac{ \pi }{ 4 })]\] \[\cos^-1 [\tan(45)]\] \[\cos^-1 [(1)]\] PUNCH IT IN A CALC and you should get 0 *please double check*